nightshade0773:

the sum of g and 381 is the same as 146 supie don't hurt my brain

KpopAsianDramaFanYT:

Look at this - https://www.mathwarehouse.com/arithmetic/multiple-calculator.php

nightshade0773:

@kpopasiandramafanyt wrote:
that just hurt my brain even more

KpopAsianDramaFanYT:

@nightshade0773 wrote:
@kpopasiandramafanyt wrote:
that just hurt my brain even more
WDYM?

nightshade0773:

@kpopasiandramafanyt wrote:
@nightshade0773 wrote:
@kpopasiandramafanyt wrote:
that just hurt my brain even more
WDYM?
i didnt even under stand it -n-

Extrinix:

"the sum of g and 381 is the same as 146" Equation created from this: g + 381 = 146 - - - - - - - - - - - - - - - - - - - - - - - - - g + 381 = 146 get g by itself - 381 -381 g = -235 your answer or 381 + g = 146 get g by itself -381 -381 g = -235 your answer, again so your answer would be `-235`.

NineTailedFox:

wow i was solving this mentally and then you replied oh well

nightshade0773:

thx every one

NineTailedFox:

i know i didnt help, sorry

nightshade0773:

its ok